OpenStudy (anonymous):

This is about a true-false test. Assume that 9 questions are answered by guessing. What is the probability of at least 6 correct answers?_______?

OpenStudy (zarkon):

compute the prob you get 6 correct then 7 ,8 then 9...add all together

OpenStudy (rulnick):

Easiest way to compute is this:

OpenStudy (rulnick):

P(6 or more correct) = P(5 or more correct) - P(exactly 5 correct)

OpenStudy (rulnick):

P(5 or more correct) = 1/2, by symmetry.

OpenStudy (rulnick):

P(exactly 5 correct) = (9 choose 5) (1/2)^9

OpenStudy (rulnick):

9 choose 5 = 9!/(5!4!) = 9(8)(7)(6)/((4)(3)(2)(1)) = 126

OpenStudy (rulnick):

So P(6 or more correct) = 1/2 - 126(1/2^9) = 1/2 - 126(1/512) = 1/2 - 126/512 = 1/2 - 63/256 = 65/126.

OpenStudy (rulnick):

Or, in decimal form, 0.25390625.

OpenStudy (zarkon):

I think you ment to type \[\frac{65}{256}\]

OpenStudy (rulnick):

(The next to last line had a typo, the 126 was supposed to be 256, obviously.)

OpenStudy (rulnick):

Yes, indeed. And the decimal value had it right.